OpenStudy (anonymous):

if n is divided by 9, the remainder is 5. what is the remainder if 3n is divided by 9?

OpenStudy (anonymous):

take the remainder, multiply by 3 and then divide by 9 and see what the reaminder is. n = 9*d + 5 => 3n = 3(9*d + 5) = 9*3d + 15 = ...

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."